OpenStudy (anonymous):

rewrite in vertex form: y= -x^2+4x-8 PLs. help I just need the guide in getting the k and H plsss.. Thanks! :)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ y= -x^2+4x-8\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ y= -x^2+4x-4-4\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ y= -(x^2-4x+4)-4\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ y= -(x-2)^2-4\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

Parabola. Opens down. Not stretched vertically or horizontally. Maximum point at (2,-4)

OpenStudy (anonymous):

thank you! but how about in fraction

OpenStudy (solomonzelman):

fraction ?

OpenStudy (solomonzelman):

I gave you the vertex form, didn't I?

OpenStudy (solomonzelman):

y = ⅕(2x²-4x+7) at first I would factor out of 2. y = ⅖(x²-2x+3½) And now, you have inside the parenthesis `2½` more than what you want to make the parenthesis a perfect square Follow this carefully. y = ⅖(x²-2x+1+2½) y = ⅖(x²-2x+1)+(2½×⅖) y = ⅖(x²-2x+1)+({5/2}×{2/5}) y = ⅖(x²-2x+1)+(1) y = ⅖(x²-2x+1)+1 y = ⅖(x-1)²+1

OpenStudy (anonymous):

Thank You so much sir! it helps a lot! :))

OpenStudy (solomonzelman):

Anytime:)